OpenStudy (anonymous):

find the inverse of f(x)= 2x^2-3

OpenStudy (anonymous):

give its verification

OpenStudy (anonymous):

@ganeshie8 pls help

OpenStudy (anonymous):

put y=2x^2-3 y+3=2x^2 y+3/2=x^2 rt(y+3/2)=x now put x in place of y and y in place of x f(x)=rt(x+3/2)

OpenStudy (anonymous):

@Faye281 is the answer ryt

OpenStudy (anonymous):

ty @Faye281

OpenStudy (anonymous):

yes but my question is how dod you get the answer the verification

OpenStudy (anonymous):

f(f^-1(x)) and f^-1(f(x))

OpenStudy (anonymous):

what is rt? sqaure root? @ajaysagar

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the verification @ajaysagar

OpenStudy (anonymous):

keep the above value in f(x) put f(rt [x+3]/2) you get f(rt [x+3]/2) =2*(x+3)/2-2 you get x

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome